OpenStudy (iwanttogotostanford):

Solve: 8(x – 2) = 5(x + 4)

OpenStudy (iwanttogotostanford):

@Lord_Frieza

OpenStudy (iwanttogotostanford):

stuck on this one too

OpenStudy (anonymous):

give me all the work you did

OpenStudy (anonymous):

then i shall give you mine so you can compare and contrast

OpenStudy (anonymous):

Let's solve your equation step-by-step. 8(x−2)=5(x+4) Step 1: Simplify both sides of the equation. 8(x−2)=5(x+4) Simplify: (Show steps) 8x−16=5x+20 Step 2: Subtract 5x from both sides. 8x−16−5x=5x+20−5x 3x−16=20 Step 3: Add 16 to both sides. 3x−16+16=20+16 3x=36 Step 4: Divide both sides by 3. 3x 3 = 36 3 x=12

OpenStudy (anonymous):

thats my work do you understand it or do you want me to further explain it?

OpenStudy (lvana):

I worked it out as @Lord_Frieza had previously stated. Hopefully seeing it all worked out helps you understand. http://puu.sh/lZ6Kv/b87106585f.png Do tell me if this helps.

OpenStudy (anonymous):

I cant click on that my school has it blocked for some reason

OpenStudy (lvana):

Hopefully this one works.~

OpenStudy (anonymous):

k that one did and it works out